OpenStudy (anonymous):

Jen has collected information about the students in her class and the month they were born. Which of the following statements would be true about this information? It is both a relation and a function. It is a relation but not a function. It is not a relation, but it is a function. It is neither a relation nor a function.

OpenStudy (cwrw238):

you can't have a student born in more than 1 month so i think its a function

OpenStudy (anonymous):

x y 150 300 250 500 360 720 480 960 f(x) =x+ 150 f(x) = 2x f(x) = f(x) = 2x- 100 © 2009, 2011 Glynlyon, Inc.

OpenStudy (cwrw238):

its also a relation so first option is correct

OpenStudy (anonymous):

Which number set correctly represents the range of the given data? {(0, 0), (2, 0), (3, 0), (5, 0)} {0} {0, 2} {0, 2, 3} {0, 2, 3, 5}

OpenStudy (cwrw238):

the second number in the paired parentheses is in the range so its {0}

OpenStudy (anonymous):

THANK YOU SO MUCH SIR!!!!!!!!!!!!!!!:)